# The Hidden Role of Anisotropies in Shaping Structure Formation in Cosmological N-Body Simulations

### v1

@@ Abstract
Initial conditions in cosmological $N$-body simulations are typically generated by displacing particles from a regular cubic lattice using a correlated field derived from the linear power spectrum, often via the Zel’dovich approximation. While this procedure reproduces the target two-point statistics (e.g., the power spectrum or correlation function), it introduces subtle anisotropies due to the underlying lattice structure. These anisotropies, invisible to angle-averaged diagnostics, become evident through directional measures such as the Angular Distribution of Pairwise Distances. Analyzing two Cold Dark Matter simulations with varying resolutions, initial redshifts, and box sizes, we show that these anisotropies are not erased but are amplified by gravitational evolution. They seed filamentary structures that persist into the linear regime, remaining visible even at redshift $z=0$. Our findings demonstrate that such features are numerical artifacts — emerging from the anisotropic coupling between the displacement field and the lattice — not genuine predictions of an isotropic cosmological model. These results underscore the importance of critically reassessing how initial conditions are constructed, particularly when probing the large-scale, quasi-linear regime of structure formation.
@@
On small scales ($r\lesssim 10,\mathrm{Mpc}/h$), non-linear gravitational collapse leads to the formation of virialized, quasi-spherical halos [11]. In this regime, the non-linear dynamics erase the memory of the initial conditions (ICs), so that the resulting structures predominantly reflect the bottom-up gravitational assembly characteristic of CDM-like structure formation, rather than preexisting features in the primordial density field.
On larger scales ($r\gtrsim 10,\mathrm{Mpc}/h$), the dynamics remain in the linear or mildly non-linear regime, where structures such as filaments can form through the coherent amplification of initial density correlations. Notably, cosmological simulations exhibit filaments that extend far beyond 10 Mpc, in some cases reaching scales comparable to the entire simulation volume — up to several hundred megaparsecs [7, 8, 9]. Despite their prominence, the physical origin of these extended structures is often not scrutinized in detail, raising important questions about their relation to ICs, finite-size effects, and numerical artifacts.
In standard CDM cosmology, particles typically travel no more than $\sim 10,\mathrm{Mpc}/h$ from their initial positions during the simulation. This implies that structures larger than this scale must emerge from correlations already present in the ICs. At early times, the two-point correlation function $\xi(r)$ of the CDM density field exhibits a clear transition: it follows a power-law decay at small scales, while at intermediate scales ($10\lesssim r\lesssim 100,\mathrm{Mpc}/h$) it approaches an exponential decay [12]. The characteristic scale of this exponential behavior — usually defined as the correlation length in statistical physics [12] — is approximately $r_{c}\approx 18\,\mathrm{Mpc}/h$ [13]. Beyond this scale, $\xi(r)$ has negligible amplitude, and no long-range correlations or coherent structures of even small amplitude are expected to be amplified under linear gravitational evolution.
@@
We argue that such large-scale filamentary structures on scales larger than 10-20 Mpc/h/h are artifacts of how the ICs are generated. In most $N$-body codes, ICs are produced by displacing particles from a regular lattice (the so-called pre-initial conditions) using a correlated Gaussian displacement field derived from the initial power spectrum, often via the Zel’dovich approximation (see, e.g., [19, 20, 21, 22] and references therein). While this procedure ensures that spherically averaged statistics, such as the power spectrum $P(k)$ or the two-point correlation function $\xi(r)$, match the theoretical model reasonably well over a finite range of spatial scales, it simultaneously introduces subtle but non-trivial anisotropies into the particle distribution. These anisotropies arise from the intrinsic spatial correlations and geometric symmetries of the regular grid used as the pre-initial configuration. Crucially, they are not fully erased by the small-amplitude, correlated displacement field applied during the generation of the ICs. These residual anisotropies are typically not captured by angle-averaged statistics. However, they become evident when analyzed using directional statistics, such as the Angular Distribution of Pairwise Distances (ADPD), which is sensitive to anisotropic clustering. The ADPD provides a more complete diagnostic of the initial density field, revealing directional features that may seed spurious structures during gravitational evolution.
@@
We emphasize that our focus is on a distinct feature — namely, anisotropic patterns in the ICs — as opposed to the issues more commonly addressed in the literature, which pertain to different kinds of discreteness effects. These effects concern how $N$-body simulations approximate the dynamics governed by the coupled Vlasov-Poisson equations in the limit of a large number of particles. Theoretically, dark matter is treated as a collisionless fluid, whose evolution is described by the Vlasov-Poisson system. In practice, however, $N$-body simulations represent this fluid using a finite set of discrete “macroparticles”whose trajectories are evolved under gravitational interactions. The fundamental assumption is that this discrete particle system provides an accurate sampling of the continuous fluid behavior. Several discreteness-related issues have been explored in the literature, including correlations induced by the initial particle sampling [23, 24], two-body relaxation effects [25, 26], and dynamical discreteness effects captured by particle linear theory [27]. (In this context, [28] proposed corrections to mitigate small-scale discreteness effects by modifying the ICs themselves.)
@@
To detect and quantify anisotropic features, we employ the ADPD as our primary tool. While the ADPD has not yet been widely adopted in cosmological analyses, it directly quantifies directional anisotropies in pairwise correlations — a statistical aspect that has been extensively studied using other approaches in the literature. For example, [29] introduced an algorithm based on persistent homology for detecting filamentary structures, where directional anisotropy is a central feature, whereas [30] analyzed pair orientations to study the alignment of galaxies within the large-scale structure. Similarly, [31] examined anisotropies in redshift-space distortions, providing foundational tools for analyzing directional correlations and [32] investigated angular correlations between galaxy spin axes, conceptually related to pairwise angular statistics. For a broader overview of techniques used to identify anisotropic features in simulations, see [33, 34, 35] who introduce various numerical methods for filament detection and the study of structure orientation.
@@ I Theoretical models
Figure 1: Upper panels: Left: power spectrum for a typical early-time Cold Dark Matter model (Eq.1). Right: corresponding two-point correlation function shown in a linear-log plot. Bottom panels: Left: correlation function displayed in a log-log scale. Right: correlation function in a linear-linear plot. Note: The amplitude is shown in arbitrary units.
The correlation function of a typical early-time CDM model is shown in Fig. 1, along with the corresponding power spectrum. The power spectrum exhibits an approximate behavior given by [13]
$P(k)\sim\frac{k}{1+\left(\frac{k}{k_{t}}\right)^{m}},$ (1)
where the turnover scale is $k_{t}\approx 0.05$ hh/Mpc, and the exponent at small scales (i.e., large $k$) is $m\approx 2.5$.
At the initial time, the correlation function — i.e., the Fourier conjugate of the power spectrum $P(k)$ — satisfies $\xi(r)<1$ for all $r$, indicating that density fluctuations are small and remain in the linear regime. The functional form of the correlation function corresponding to a power spectrum of the type given in Eq. 1 can be computed numerically (see Fig.1). In summary it behaves as follows: within the range $\sim 10$$100\,\mathrm{Mpc}$/$h$, it exhibits an exponential decay of the form
$\xi(r)\sim\exp\left(-\frac{r}{r_{c}}\right),$ (2)
@@
Physically, the meaning can be clarified by considering a general classification of stochastic processes in terms of their spectral properties. In a general classification of stochastic processes, consider the case where the power spectrum behaves as $P(k)\sim k^{n}$ in the limit $k\rightarrow 0$. The corresponding mass variance, defined as
$\sigma^{2}(r)=\frac{\langle\Delta N\rangle^{2}}{\langle N\rangle^{2}}\sim r^{-\gamma},$
obeys the relation $\gamma=3+n$ for 3<n<1-3<n<1, and $\gamma=4$ for $n>1$. The random (Poisson) case corresponds to a power spectrum with $n=0$, i.e., $P(k)\sim\text{const.}$, which leads to a mass variance that scales as $\sigma^{2}(r)\sim r^{-3}$. In contrast, the Harrison-Zel’dovich spectrum — characteristic of the small-$k$ behavior of CDM-like models — corresponds to $n=1$, for which the mass variance behaves as $\sigma^{2}(r)\sim\log(r)\,r^{-4}$.
From a physical point of view, mass fluctuations in a super-homogeneous system scale with the surface area of the sampling region, rather than its volume. This is in stark contrast to the Poisson case, where fluctuations scale with volume. The suppressed large-scale fluctuations in CDM-like models reflect the long-range order and enhanced uniformity characteristic of super-homogeneous fields. Thus, the correlation properties of the CDM theoretical model at early times can be summarized as follows:
@@
    For $r<10,\mathrm{Mpc}/h$, there are positive, but small in amplitude, power-law correlations: $P(k)\sim k^{-2}$ and $\xi(r)\sim r^{-1}$. In this regime, during time evolution up to the present epoch non-linear gravitational clustering distorts the initial shape of $\xi(r)$ when its amplitude becomes of order unity, leading to a change in the power-law index.
@@
    For $r\gtrsim 120\,\mathrm{Mpc}/h$, where the power spectrum behaves as $P(k)\sim k$, the correlation function follows $\xi(r)\sim-r^{-4}$, indicating an anti-correlated regime and, consequently, the absence of large-scale structure. This range of scales also remain in the linear regime during the whole evolution.
@@
This length scale sets the size of a typical void, i.e., a region devoid of particles. It corresponds to the scale of non-linearity: for $r>\lambda$, the average density is well defined and fluctuations are small. Therefore, for $r>\lambda$, the correlation properties reflect those of the initial conditions, modulated only by a linear amplification due to gravitational clustering in an expanding universe.
@@ II.1 The Zel’dovich approximation
Assuming statistical homogeneity and isotropy of the displacement field $\vec{u}(\vec{r})$, one can transform Eq. (5) into Fourier space. Taking the expectation value of the square modulus of the density fluctuations, i.e., the power spectrum (PS), one obtains:
@@
$\lim_{k\rightarrow 0}P(k)\sim k^{n}\quad\text{with}\quad n>-1\;.$
This implies that by applying a suitably correlated displacement field to a uniform density background, one can construct a fluctuation field with a power spectrum of the form $P(k)\sim k^{n}$ for $n\leq-1$.
@@ II.2 Pre-initial conditions
As discussed in the introduction, a number of studies have investigated the so-called “discreteness effects,” [38, 25, 39, 27, 26, 23, 28]. In this work, we focus on a related but distinct issue: whether, from the perspective of statistical characterization, the ICs provide a fair realization of the underlying theoretical model. We show that the answer to this problem depends critically on the statistical diagnostics employed. While standard measures such as the power spectrum and two-point correlation function offer angle-averaged information, they may fail to capture directional features that violate statistical isotropy. Our aim is to assess these subtle anisotropies using more sensitive directional statistics.
@@
As we will discuss below, the power spectrum of ICs generated from a regular cubic grid typically agrees well with the theoretical prediction for wavenumbers smaller than the Nyquist frequency, $k_{N}=\pi n_{0}^{1/3}$, where $n_{0}$ is the average number density. However, it is important to emphasize that the power-spectrum of a perfect 3D cubic lattice is non-zero only at specific discrete wave-vectors (including some below the Nyquist frequency) and zero elsewhere. As a result, for certain wave-vectors, the agreement with the theoretical spectrum is not well satisfied. This discrepancy is often overlooked because the power-spectrum is computed through spherical averaging, which implicitly assumes statistical isotropy. In contrast, a cubic lattice possesses only discrete translational invariance and exhibits strong anisotropies in Fourier space [12]. Therefore, spherical averaging can obscure underlying directional artifacts that are present in the ICs due to the symmetry of the lattice.
@@
However, it is far from clear that using a cubic lattice or glass configurations as a pre-initial condition entirely resolve this issue. Both represent unstable point configurations under gravity, and even a small applied perturbation may result in an evolution that depends sensitively on the residual correlations or fluctuations present in the unperturbed configuration.
In principle, such discreteness effects can be minimized by refining the discretization, — i.e., increasing the number of particles to better sample the underlying density field. In practice, however, $N$-body simulations often probe the non-linear regime of structure formation at scales comparable to the discretization length. Under these conditions, the influence of pre-initial correlations becomes, at the very least, problematic [40, 41, 15].
As previously noted, a lattice is not an isotropic system: intrinsic anisotropies of the lattice persist in the particle distribution even after the displacement field has been applied. The question we aim to explore below is whether these residual anisotropies have any dynamical effects during the subsequent gravitational evolution.
Note that an alternative method for generating the pre-initial particle distribution for an N-body simulation consists in the construction of a glass-like configuration [42]: this is generated by starting from a Poisson distribution and running an N-body code with the sign of the acceleration reversed (see [22] and reference therein). The main motivation for using a glass configuration instead of a grid is to eliminate the symmetries inherent to a lattice. The resulting distribution is highly isotropic but still exhibits long-range correlations akin to those of a lattice [15]. In this work, we do not consider the case of glass-like pre-initial conditions; a more detailed study of this scenario will be presented in a forthcoming paper.

### v2

@@ Abstract
Initial conditions in cosmological $N$-body simulations are typically generated by displacing particles from a regular cubic lattice using a correlated field derived from the linear power spectrum, often via the Zel’dovich approximation. While this procedure reproduces the target two-point statistics (e.g., the power spectrum or correlation function), it introduces subtle anisotropies due to the underlying lattice structure. These anisotropies, invisible to angle-averaged diagnostics, become evident through directional measures such as the Angular Distribution of Pairwise Distances. Analyzing two Cold Dark Matter simulations with varying resolutions and box sizes, we show that these anisotropies are not erased but are amplified by gravitational evolution. They seed filamentary structures that persist into the linear regime, remaining visible even at redshift $z=0$. Our findings demonstrate that such features are numerical artifacts — emerging from the anisotropic coupling between the displacement field and the lattice — not genuine predictions of an isotropic cosmological model. These results underscore the importance of critically reassessing how initial conditions are constructed, particularly when probing the large-scale, quasi-linear regime of structure formation.
Keywords: Cosmology; Fluctuation phenomena random processes noise and Brownian motion
@@
On small scales ($r\lesssim 10\,\mathrm{Mpc}/h$), non-linear gravitational collapse leads to the formation of virialized, quasi-spherical halos [11]. In this regime, the non-linear dynamics erase the memory of the initial conditions (ICs), so that the resulting structures predominantly reflect the bottom-up gravitational assembly characteristic of CDM-like structure formation, rather than preexisting features in the primordial density field.
On larger scales ($r\gtrsim 10\,\mathrm{Mpc}/h$), the dynamics remain in the linear or mildly non-linear regime, where structures such as filaments can form through the coherent amplification of initial density correlations. Notably, cosmological simulations exhibit filaments that extend far beyond 10 Mpc, in some cases reaching scales comparable to the entire simulation volume — up to several hundred megaparsecs [7, 8, 9]. Despite their prominence, the physical origin of these extended structures is often not scrutinized in detail, raising important questions about their relation to ICs, finite-size effects, and numerical artifacts.
In standard CDM cosmology, particles typically travel no more than $\sim 10\,\mathrm{Mpc}/h$ from their initial positions during the simulation. This implies that structures larger than this scale must emerge from correlations already present in the ICs. At early times, the two-point correlation function $\xi(r)$ of the CDM density field exhibits a clear transition: it follows a power-law decay at small scales, while at intermediate scales ($10\lesssim r\lesssim 100\,\mathrm{Mpc}/h$) it approaches an exponential decay [12]. The characteristic scale of this exponential behavior — usually defined as the correlation length in statistical physics [12] — is approximately $r_{c}\approx 18\,\mathrm{Mpc}/h$ [13]. Beyond this scale, $\xi(r)$ has negligible amplitude, and no long-range correlations or coherent structures of even small amplitude are expected to be amplified under linear gravitational evolution.
@@
We argue that such large-scale filamentary structures on scales larger than 10-20 Mpc/h/h are artifacts of how the ICs are generated. In most $N$-body codes, ICs are produced by displacing particles from a regular lattice (the so-called pre-ICs) using a correlated Gaussian displacement field derived from the initial power spectrum (PS), often via the Zel’dovich approximation (see, e.g., [19, 20, 21, 22] and references therein). While this procedure ensures that spherically averaged statistics, such as the PS $P(k)$ or the two-point correlation function $\xi(r)$, match the theoretical model reasonably well over a finite range of spatial scales, it simultaneously introduces subtle but non-trivial anisotropies into the particle distribution. These anisotropies arise from the intrinsic spatial correlations and geometric symmetries of the regular grid used as the pre-initial configuration. Crucially, they are not fully erased by the small-amplitude, correlated displacement field applied during the generation of the ICs. These residual anisotropies are typically not captured by angle-averaged statistics. However, they become evident when analyzed using directional statistics, such as the Angular Distribution of Pairwise Distances (ADPD), which is sensitive to anisotropic clustering. The ADPD provides a more complete diagnostic of the initial density field, revealing directional features that may seed spurious structures during gravitational evolution.
@@
We emphasize that our focus is on a distinct feature — namely, anisotropic patterns in the ICs — as opposed to the issues more commonly addressed in the literature, which pertain to different kinds of discreteness effects. These effects concern how $N$-body simulations approximate the dynamics governed by the coupled Vlasov-Poisson equations in the limit of a large number of particles. Theoretically, dark matter is treated as a collisionless fluid, whose evolution is described by the Vlasov-Poisson system. In practice, however, $N$-body simulations represent this fluid using a finite set of discrete “macroparticles”whose trajectories are evolved under gravitational interactions. The fundamental assumption is that this discrete particle system provides an accurate sampling of the continuous fluid behavior. Several discreteness-related issues have been explored in the literature, including correlations induced by the initial particle sampling [23, 24], impact of particle noise on cosmological N-body simulations [25], two-body relaxation effects [26, 27], and dynamical discreteness effects captured by particle linear theory [28]. (In this context, [29] proposed corrections to mitigate small-scale discreteness effects by modifying the ICs themselves.)
The impact of artificial anisotropies introduced by pre-ICs has been systematically investigated in several recent studies. [37] quantified these effects by comparing cosmological simulations initialized from a regular grid with those based on a glass configuration. Their results demonstrated that glass-based pre-ICs yield more stable and accurate outcomes across a wide range of scales, consistently outperforming the standard grid-based approach.
Building on this, [38] examined the anisotropy inherent in gravitational force calculations arising from the 3-torus topology imposed by periodic boundary conditions. To isolate this effect, they employed glass-like ICs, thereby eliminating grid-related artifacts. Their analysis suggests that the initial grid not only seeds small-scale preferential directions but may also amplify the intrinsic anisotropic features of the periodic gravitational field.
More recently, [39] studied the residual anisotropy sourced by cubic lattice pre-ICs and showed that such anisotropies are not fully erased by nonlinear gravitational evolution. While the amplitude of the effect remains modest, the anisotropy signal was found to be statistically significant in the orientations of haloes and filaments. Taken together, these studies demonstrate that lattice-induced anisotropies can leave measurable imprints on the large-scale structure, even at late times. Moreover, they collectively support the conclusion that glass-based pre-ICs are more effective in mitigating such artificial anisotropies, reinforcing the case for their adoption in high-precision cosmological simulations.
@@
In this work, we employ the ADPD as our primary tool to detect and quantify anisotropic features and we demonstrate that the anisotropy induced by cubic lattice pre-initial conditions persists over time and plays a role in the formation of large-scale structures. While the ADPD has not yet been widely adopted in cosmological analyses, it directly quantifies directional anisotropies in pairwise correlations — a statistical aspect that has been extensively studied using other approaches in the literature. For example, [30] introduced an algorithm based on persistent homology for detecting filamentary structures, where directional anisotropy is a central feature, whereas [31] analyzed pair orientations to study the alignment of galaxies within the large-scale structure. Similarly, [32] examined anisotropies in redshift-space distortions, providing foundational tools for analyzing directional correlations and [33] investigated angular correlations between galaxy spin axes, conceptually related to pairwise angular statistics. For a broader overview of techniques used to identify anisotropic features in simulations, see [34, 35, 36] who introduce various numerical methods for filament detection and the study of structure orientation.
@@ I Theoretical models
Figure 1: Power spectrum (left) and corresponding two-point correlation function in linear-log (middle) and log-log (right) plots for a typical early-time Cold Dark Matter model (see Eq. 1). The vertical red line marks the zero-crossing scale $r_{t}\approx 120\mathrm{Mpc}$/$h$. Amplitudes are shown in arbitrary units.
The correlation function of a typical early-time CDM model is shown in Fig. 1, along with the corresponding PS. The PS exhibits an approximate behavior given by [13]
$P(k)=A\frac{k}{1+\left(\frac{k}{k_{t}}\right)^{m}},$ (1)
where $A$ is a normalization constant, the turnover scale is $k_{t}\approx 0.05$ hh/Mpc, and the exponent at small scales (i.e., large $k$) is $m\approx 2.5$. For the sake of clarity, we have adopted a PS without baryon acoustic oscillations (BAO), as these features are not essential to the main arguments of our work. Including BAO would merely introduce small-amplitude wiggles in the correlation function at scales around $100$ Mpc/h/h, without affecting our conclusions.
At the initial time, the correlation function — i.e., the Fourier conjugate of the PS $P(k)$ — satisfies $\xi(r)<1$ for all $r$, indicating that density fluctuations are small and remain in the linear regime. The functional form of the correlation function corresponding to a PS of the type given in Eq. 1 can be computed numerically (see Fig.1). In summary it behaves as follows: within the range $\sim 10$$100\,\mathrm{Mpc}$/$h$, it exhibits an exponential decay of the form
$\xi(r)\propto\exp\left(-\frac{r}{r_{c}}\right),$ (2)
@@
Physically, the meaning of the concept of super-homogeneous can be clarified by considering a general classification of stochastic processes in terms of their spectral properties. In a general classification of stochastic processes, consider the case where the PS behaves as $P(k)\propto k^{n}$ in the limit $k\rightarrow 0$. The corresponding mass variance, defined as
$\sigma^{2}(r)=\frac{\langle\Delta N\rangle^{2}}{\langle N\rangle^{2}}\propto r^{-\gamma},$
obeys the relation $\gamma=3+n$ for 3<n<1-3<n<1, and $\gamma=4$ for $n>1$. The random (Poisson) case corresponds to a PS with $n=0$, i.e., $P(k)\propto\text{const.}$, which leads to a mass variance that scales as $\sigma^{2}(r)\propto r^{-3}$. In contrast, the Harrison-Zel’dovich spectrum — characteristic of the small-$k$ behavior of CDM-like models — corresponds to $n=1$, for which the mass variance behaves as $\sigma^{2}(r)\propto\log(r)\,r^{-4}$.
This behavior corresponds to the fact that mass fluctuations in a super-homogeneous system scale with the surface area of the sampling region, rather than its volume. This is in stark contrast to the Poisson case, where fluctuations scale with volume. The suppressed large-scale fluctuations in CDM-like models reflect the long-range order and enhanced uniformity characteristic of super-homogeneous fields. Thus, the correlation properties of the CDM theoretical model at early times can be summarized as follows:
@@
    For $r<10,\mathrm{Mpc}/h$, there are positive, but small in amplitude, power-law correlations: $P(k)\propto k^{-2}$ and $\xi(r)\propto r^{-1}$. In this regime, during time evolution up to the present epoch non-linear gravitational clustering distorts the initial shape of $\xi(r)$ when its amplitude becomes of order unity, leading to a change in the power-law index.
@@
    For $r\gtrsim 120\,\mathrm{Mpc}/h$, where the PS behaves as $P(k)\propto k$, the correlation function follows $\xi(r)\propto-r^{-4}$, indicating an anti-correlated regime and, consequently, the absence of large-scale structure. This range of scales also remain in the linear regime during the whole evolution.
@@
This length scale sets the size of a typical void, i.e., a region devoid of particles. It corresponds to the scale of non-linearity: for $r>\lambda$, the average density is well defined and fluctuations are small. Therefore, for $r>\lambda$, the correlation properties reflect those of the ICs, modulated only by a linear amplification due to gravitational clustering in an expanding universe.
@@ II.1 The Zel’dovich approximation
Assuming statistical homogeneity and isotropy of the displacement field $\vec{u}(\vec{r})$, one can transform Eq. (5) into Fourier space. Taking the expectation value of the square modulus of the density fluctuations, i.e., the PS, one obtains:
@@
$\lim_{k\rightarrow 0}P(k)\propto k^{n}\quad\text{with}\quad n>-1\;.$
This implies that by applying a suitably correlated displacement field to a uniform density background, one can construct a fluctuation field with a PS of the form $P(k)\propto k^{n}$ for $n\leq-1$.
@@ II.2 Pre-initial conditions
As discussed in the introduction, a number of studies have investigated the so-called “discreteness effects” [42, 26, 43, 28, 27, 23, 25, 29, 37, 38, 39]. In this work, we focus on a related but distinct issue: whether, from the perspective of statistical characterization, the ICs provide a fair realization of the underlying theoretical model. We show that the answer to this problem depends critically on the statistical diagnostics employed. While standard measures such as the PS and two-point correlation function offer angle-averaged information, they may fail to capture directional features that violate statistical isotropy. Our aim is to assess these subtle anisotropies using more sensitive directional statistics.
@@
As we will discuss below, the PS of ICs generated from a regular cubic grid typically agrees well with the theoretical prediction for wave-numbers smaller than the Nyquist frequency, $k_{N}=\pi n_{0}^{1/3}$, where $n_{0}$ is the average number density. However, it is important to emphasize that the PS of a perfect 3D cubic lattice is non-zero only at specific discrete wave-vectors (including some below the Nyquist frequency) and zero elsewhere. As a result, for certain wave-vectors, the agreement with the theoretical spectrum is not well satisfied. This discrepancy is often overlooked because the PS is computed through spherical averaging, which implicitly assumes statistical isotropy. In contrast, a cubic lattice possesses only discrete translational invariance and exhibits strong anisotropies in Fourier space [12]. Therefore, spherical averaging can obscure underlying directional artifacts that are present in the ICs due to the symmetry of the lattice.
@@
However, it is far from clear that using a cubic lattice or glass configurations as a pre-ICs entirely resolve this issue. Both represent unstable point configurations under gravity, and even a small applied perturbation may result in an evolution that depends sensitively on the residual correlations or fluctuations present in the unperturbed configuration. In principle, such discreteness effects can be minimized by refining the discretization, — i.e., increasing the number of particles to better sample the underlying density field. In practice, however, $N$-body simulations often probe the non-linear regime of structure formation at scales comparable to the discretization length. Under these conditions, the influence of pre-initial correlations becomes, at the very least, problematic [44, 45, 15].
As previously noted, a lattice is not an isotropic system: intrinsic anisotropies of the lattice persist in the particle distribution even after the displacement field has been applied [37, 38, 39]. The question we aim to explore below is whether these residual anisotropies have any dynamical effects during the subsequent gravitational evolution.
Note that an alternative method for generating the pre-initial particle distribution for an N-body simulation consists in the construction of a glass-like configuration [46]: this is generated by starting from a Poisson distribution and running an N-body code with the sign of the acceleration reversed (see [22] and reference therein). The main motivation for using a glass configuration instead of a grid is to eliminate the symmetries inherent to a lattice. The resulting distribution is highly isotropic but still exhibits long-range correlations akin to those of a lattice [15]. In this work, we do not consider the case of glass-like pre-ICs; a more detailed study of this scenario will be presented in a forthcoming paper.